Fermi National Laboratory
 

Operation Summary for November 4, 2009 - November 6, 2009

 

Wednesday Nov. 4

The day shift began with Operations monitoring the Tevatron (TeV) store 7321 (86.78E30) and the Recycler stash (228.5E10), with the Antiproton Source stacking (21.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7321 at 12:26 PM. Just before the store was terminated, cryo system personnel from CHL adjusted the TeV sector A2 high-energy lead flow. Operations put the TeV into shot setup and established store 7322 at 2 PM with an average initial luminosity (AIL) of 230E30. Linac, Booster, NUMI, and MI made accesses for necessary repairs. The Booster, NUMI, and MI accesses lasted through the end of the shift.

During the evening shift, Operators had the NuMI power supplies on and ramping by 5:30 PM. Operations had Booster up and running with beam to the Booster dump by 6:20 PM. MiniBooNE operations resumed by 6:38 PM. MI was on and ramping by 6:50 PM. Operations established beam to Pbar and NuMI by 7:12 PM. Operations reported that a MI RF station (MIRF4) would not stay running, so they bypassed the station. Operators accessed MI-31 to restart the Pelletron BPM node.

Thursday Nov. 5

The midnight shift began with Operations monitoring store 7322 (82.43E30) and the stash (109.4E10), with Pbar stacking (15.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that here was some downtime when the MI south anode power supply tripped off. The MI vacuum system front end (MIVAC) required a reboot.

During the day shift, Operations terminated store 7322 at 9:45 PM. Shot setup followed, but was delayed due to MI RF trips on a watchdog alarm and by MI SBD not responding. Operations established store 7323 at noon with an AIL of 239.6E30. Also at noon, RF personnel brought MIRF4 back online. Recycler personnel conducted studies after the antiprotons were off-loaded. Store 7323 aborted at 2:52 PM when the TeV sector F3 safety coordinator failed. The Pbar lithium lens tripped off due to an external interlock glitch; this caused some downtime. Linac personnel tested the Neutron Therapy Facility (NTF) 58-degree magnet power supply for most of the shift. The Booster chopper (B:CHOP) began causing problems, but Booster personnel made a temporary repair to keep it running until they can replace the “ON” tube. Cryo system personnel replaced the TeV sector A48 lead flow solenoid. Mechanical personnel filled the TeV sector B48 separators with fluorinert.

During the evening shift, EE Support personnel had the TeV safety coordinator changed out by 5 PM. Operations turned on the TeV and conducted orbit checks until 7 PM. Operations put the TeV into shot setup and established store 7325 at 6:30 PM with an AIL of 146.7E30. The lithium lens tripped off a couple of times on an external interlock glitch - target system personnel helped operators reset the trip. MIVAC tripped once this shift.

Friday Nov. 6

The midnight shift began with Operations monitoring store 7325 (80.43E30) and the stash (111.7E10), with Pbar stacking (11.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Linac RF station (LRF3) tripped off with a Power Amplifier (PA) LCW leak indication (the leak was actually on the driver tube - the PA and driver leak interlocks are summed together). Operations called the Linac on-call technician; he brought the station back online. Operations reported that the MI RF system started tripping off due to stations 1 and 5 when Linac came back up. Operations had to lower the beam intensity for stacking and NuMI to stop the trips. The Pbar lithium lens tripped off several times with the same indications as on the last two shifts; Operations followed the instructions left by the Pbar personnel and reset the lens as required. MIVAC tripped once this shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Linac personnel to repair LRF3, Booster personnel to repair the chopper, and power supply personnel to repair the MI-10 lower power supply as soon as NTF finishes treating their last patient later this morning.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/06/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory